OpenStudy (gabylovesyou):

how do i graph f(x) = (9x^2 - 36) / (3x + 6) @misty1212

imqwerty (imqwerty):

ok so we have this->\(\huge\ f(x)= \frac{ 9x^2-36 }{ 3x+6 }\) try to apply this identity in the numerator->\(\large\color{violet}{ a^2-b^2=(a+b)(a-b)}\)

OpenStudy (anonymous):

you end up graphing \(y=3x-6\) a line

OpenStudy (gabylovesyou):

9(x+2)(x-2)

imqwerty (imqwerty):

what did you do

OpenStudy (gabylovesyou):

difference of squares

imqwerty (imqwerty):

ok thats not correct see it like this->\(\huge\ f(x)= \frac{ (3x)^2-(6)^2 }{ 3x+6 }\) now try :)

OpenStudy (gabylovesyou):

factor out using GCF?

imqwerty (imqwerty):

no 1st you apply this identity in the numerator and factor the terms-\(\large\color{violet}{ a^2-b^2=(a+b)(a-b)}\) and then you cancel out the common factors in numerator and denominator

OpenStudy (anonymous):

it is the same thing, makes no difference

imqwerty (imqwerty):

oh sorry i didn't noticed >.<

OpenStudy (anonymous):

you can factor as \(9(x^2-4)=9(x+2)(x-2)\)

OpenStudy (gabylovesyou):

yea thats what i got.. and then ?

OpenStudy (anonymous):

and denominator as \[3(x+2)\]

OpenStudy (anonymous):

then cancel and get \[3(x-2)\]

OpenStudy (anonymous):

or \[y=3x-6\] same thing either way

OpenStudy (gabylovesyou):

ok and what number will be circled when graphing..

OpenStudy (anonymous):

if \(x+2=0\) then \(x=-2\)

OpenStudy (anonymous):

so you can't put in \(-2\) that will have the circle

OpenStudy (anonymous):

second coordinate will be \[3(-2)-6=-12\]

OpenStudy (gabylovesyou):

these are my choices.. i know its not A or B..

OpenStudy (anonymous):

go with C as usual

OpenStudy (anonymous):

the one with the circle at \((-2,-12)\)

OpenStudy (gabylovesyou):

ok but how come ?

OpenStudy (anonymous):

you cancelled the \(x+2\) from the denominator , which is fine, but the domain does not include \(-2\) since the denominator would be zero there

OpenStudy (anonymous):

if you put \(-2\) in the cancelled expression \(3x-6\) you get \(-12\)

OpenStudy (gabylovesyou):

oh....... okok let me do another one to make sure.. ill tag you : )